# ELITR-Bench: A Meeting Assistant Benchmark for Long-Context Language Models

### v1

@@ 1 Introduction
The context window of Large Language Models (LLMs) has recently undergone a significant expansion, scaling from a few thousand tokens to tens or even hundreds of thousands [6, 40, 24, 7, 3]. As a consequence, benchmarks have emerged to assess LLMs’ long-range abilities, tackling the specific challenges of Question Answering (QA) on long documents [1, 22, 2, 23, 27, 43]. However, while previous datasets focusing on long-context models offer longitudinal evaluations across different tasks, they often provide only superficial analyses of each task. The covered tasks are also often generic – e.g., questions on Wikipedia [23] – and thus not particularly suitable for realistic, focused applications.
In contrast, our work advocates for a situated evaluation of long-context LLM performance within specific, real-world scenarios. As a practical illustration, consider a meeting assistant that allows users to inquire about meetings they did not attend. Given that hour-long meeting transcripts must fit within the agent’s context window, proficient handling of long contexts is a prerequisite. This paper then introduces the first benchmark – to the best of our knowledge – for evaluating long-context LLMs on a realistic meeting assistant task. Our benchmark, named ELITR-Bench, [^1] We release the data for our benchmark at https://github.com/utter-project/UTTER-MS9-meetingdata/tree/master/ELITR-Bench. is built upon the meeting transcripts of the past ELITR project [29]. These transcripts have been obtained by Automatic Speech Recognition (ASR) with minimal human corrections – yielding long, noisy documents of oral nature that present unique challenges for LLMs. Our extensive experiments on ELITR-Bench with 9 recent long-context LLMs showed a gap between proprietary and open-source models that is emphasized when questions are asked sequentially within a conversation rather in a QA mode. We also provide an in-depth validation of our GPT-4-based evaluation through a crowdsourcing study, showing a high correlation with human judges’ scores. However, we also point out GPT-4’s limited ability to distinguish beyond three score levels.
@@ 2 Related work
Recent works explored the use of LLMs such as GPT-4 as judges to evaluate responses on open-ended questions. Zheng et al. [44] measured agreement between LLM and human evaluators while introducing two datasets (MT-bench and Chatbot Arena). They showed that LLM judges like GPT-4 can match both controlled and crowdsourced human annotations, achieving over 80% agreement – the same level of agreement between humans. He et al. [15] evaluated the performance of GPT-4 against 415 crowdsourcing human labelers. Despite employing best quality control practices, the highest labeling accuracy achieved through crowdsourcing was 81.5% whereas GPT-4 obtained 83.6%. As in certain scenarios, employing proprietary LLMs as evaluators can pose challenges due to their closed-source nature, Kim et al. [18] introduced Prometheus, an open-source LLM fine-tuned for evaluation. In this work, we compare LLMs-as-a-judge (GPT-4 and Prometheus) with expert and crowdsourcing-based human evaluators to assess responses generated by several long-context models on ELITR-Bench.
@@ 5.1 Main results
Looking at the three settings in Table 2, we observe that GPT-4 clearly dominates over all other approaches with an average score that is always above 8.888While one might argue that GPT-4 is unfairly advantaged due to the use of a GPT-4 judge, we show in Section 6.2 that the dominance of this model is observed for other evaluators as well. GPT-3.5 obtained a slightly lower average score  – around 7 – that still outperforms open-source LLMs. Among these, differences are smaller with scores close to 6 on the single-turn setting, and ranging from 4 to 6 on the multi-turn settings. Nonetheless, we can note that Vicuna-13B-v1.5 is the open-source approach that performed the most favorably overall on the three settings. Interestingly, the results in the single-turn and multi-turn modes show large discrepancies for open-source models – even when the question set is exactly the same, for ELITR-Bench-QA. This seems to indicate that open-source long-context LLMs get distracted by the previous questions and answers, which affects their performance. In contrast, GPT-4 is instead able to increase its performance between the single-turn mode and the multi-turn mode. Comparing the results of the QA and Conv settings in the multi-turn mode, we found only minimal differences. This can be explained by the small number of questions that differ between QA and Conv (16 for the dev set and 17 for the test set). In Section 5.3, we analyze the results on this subset of differentiating questions to better understand the impact of the benchmark setting (QA vs Conv) and inference mode (single-turn vs multi-turn).
@@ 6.1 Compared evaluators
  • Prometheus [18]: This fine-tuned model was originally proposed to provide an open-source alternative to using GPT-4 for score rubric-based evaluation. We used the Prometheus-13B-v1.0999https://huggingface.co/kaist-ai/prometheus-13b-v1.0 model, with a prompt similar to the one adopted for GPT-4 – the only difference is that the score rubric is re-scaled to a 1-5 range to fit Prometheus’ expected format and multiplied by 2 in post-processing to be comparable to other scores. The prompt is available in Appendix E (Figs. 7 and 8).
@@ Acknowledgments
The complete set of ELITR-Bench (question, ground-truth answer) pairs, along with metadata, is provided in JSON format at https://github.com/utter-project/UTTER-MS9-meetingdata/tree/master/ELITR-Bench. We also indicate for each question the responses generated by the different long-context LLMs considered in this paper, as well as the evaluation score attributed by the GPT-4 judge and other studied evaluators.
@@ A.1 Compared models and hardware details
The inference was done on a single A100 GPU with 80GB memory. In preliminary experiments, we also attempted to include the Mistral-7B-Instruct-v0.2111111https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model in our study, as this model supports a context of up to 32K tokens. However, running this model on ELITR-Bench led to a GPU out-of-memory error on the A100, and thus we discarded it.
@@ A.2 Configuration search on ELITR-Bench-QA’s dev set
The cost of the two-step configuration search amounted to approximately \$150.141414We assessed the cost of performing the evaluation of a single model on the 141 dev set questions to \$3 approximately. As we evaluated 7 models on 6 configurations in the first step, and 7 models on 1 configuration in the second step, this yields \$147. To limit excessive expenses, we used the same model configuration for the different settings we experimented in (single-turn ELITR-Bench-QA, multi-turn ELITR-Bench-QA, and multi-turn ELITR-Bench-Conv).

### v2

@@ 1 Introduction
The context window of Large Language Models (LLMs) has recently undergone a significant expansion, scaling from a few thousand tokens to tens or even hundreds of thousands [8, 45, 28, 9, 4]. As a consequence, benchmarks have emerged to assess LLMs’ long-range abilities, tackling the specific challenges of Question Answering (QA) on long documents [1, 25, 3, 26, 31, 48]. However, while previous datasets focusing on long-context models offer longitudinal evaluations across different tasks, they often provide only superficial analyses of each task. The covered tasks are also often generic – e.g., questions on Wikipedia [26] – and thus not particularly suitable for realistic, focused applications. [^1] Aside from this very recent work [27] (WildBench), which offers an automated evaluation framework for assessing LLMs with complex real-world user queries.
In contrast, our work advocates for a situated evaluation of long-context LLM performance within specific, real-world scenarios. As a practical illustration, consider a meeting assistant that allows users to inquire about meetings they did not attend. Given that hour-long meeting transcripts must fit within the agent’s context window, proficient handling of long contexts is a prerequisite. This paper then introduces the first benchmark – to the best of our knowledge – for evaluating long-context LLMs on a realistic meeting assistant task. Our benchmark, named ELITR-Bench, [^2] We release the data associated to our benchmark, as well as the generations produced by compared models and their evaluation, at https://github.com/utter-project/ELITR-Bench. is built upon the meeting transcripts of the past ELITR project [33]. These transcripts have been obtained by Automatic Speech Recognition (ASR) with minimal human corrections – yielding long, noisy documents of oral nature that present unique challenges for LLMs. Our extensive experiments on ELITR-Bench with 9 recent long-context LLMs showed a gap between proprietary and open-source models that is emphasized when questions are asked sequentially within a conversation rather in a QA mode. We also provide an in-depth validation of our GPT-4-based evaluation through a crowdsourcing study, showing a high correlation with human judges’ scores. However, we also point out GPT-4’s limited ability to distinguish beyond three score levels.
@@ 2 Related work
Recent works explored the use of LLMs such as GPT-4 as judges to evaluate responses on open-ended questions. Zheng et al. [49] measured agreement between LLM and human evaluators while introducing two datasets (MT-bench and Chatbot Arena). They showed that LLM judges like GPT-4 can match both controlled and crowdsourced human annotations, achieving over 80% agreement – the same level of agreement between humans. He et al. [17] evaluated the performance of GPT-4 against 415 crowdsourcing human labelers. Despite employing best quality control practices, the highest labeling accuracy achieved through crowdsourcing was 81.5% whereas GPT-4 obtained 83.6%. As in certain scenarios, employing proprietary LLMs as evaluators can pose challenges due to their closed-source nature, Kim et al. [21] introduced Prometheus, an open-source LLM fine-tuned for evaluation. Recently, Bavaresco et al. [5] introduced Judge-Bench, a collection of 20 NLP datasets with human annotations for evaluating LLMs’ ability to replicate human judgments. In this work, we compare LLMs-as-a-judge (GPT-4 and Prometheus) with expert and crowdsourcing-based human evaluators to assess responses generated by several long-context models on ELITR-Bench.
@@ 5.1 Main results
Looking at the three settings in Table 2, we observe that GPT-4 clearly dominates over all other approaches with an average score that is always above 8.101010While one might argue that GPT-4 is unfairly advantaged due to the use of a GPT-4 judge, we show in Section 6.2 that the dominance of this model is observed for other evaluators as well. GPT-3.5 obtained a slightly lower average score  – around 7 – that still outperforms open-source LLMs. Among these, differences are smaller with scores close to 6 on the single-turn setting, and ranging from 4 to 6 on the multi-turn settings. Nonetheless, we can note that Vicuna-13B-v1.5 is the open-source approach that performed the most favorably overall on the three settings. Interestingly, the results in the single-turn and multi-turn modes show large discrepancies for open-source models – even when the question set is exactly the same, for ELITR-Bench-QA. This seems to indicate that open-source long-context LLMs get distracted by the previous questions and answers, which affects their performance. In contrast, GPT-4 is instead able to increase its performance between the single-turn mode and the multi-turn mode. Comparing the results of the QA and Conv settings in the multi-turn mode, we found only minimal differences. This can be explained by the small number of questions that differ between QA and Conv (16 for the dev set and 17 for the test set). In Section 5.3, we analyze the results on this subset of differentiating questions to better understand the impact of the benchmark setting (QA vs Conv) and inference mode (single-turn vs multi-turn).
@@ 6.1 Compared evaluators
  • Prometheus [21]: This fine-tuned model was originally proposed to provide an open-source alternative to using GPT-4 for score rubric-based evaluation. We used the Prometheus-13B-v1.0111111https://huggingface.co/kaist-ai/prometheus-13b-v1.0 model, with a prompt similar to the one adopted for GPT-4 – the only difference is that the score rubric is re-scaled to a 1-5 range to fit Prometheus’ expected format and multiplied by 2 in post-processing to be comparable to other scores. The prompt is available in Appendix E (Figs. 7 and 8).
@@ Acknowledgments
The complete set of ELITR-Bench (question, ground-truth answer) pairs, along with metadata, is provided in JSON format at https://github.com/utter-project/ELITR-Bench. We also indicate for each question the responses generated by the different long-context LLMs considered in this paper, as well as the evaluation score attributed by the GPT-4 judge and other studied evaluators.
@@ A.1 Compared models and hardware details
The inference was done on a single A100 GPU with 80GB memory. In preliminary experiments, we also attempted to include the Mistral-7B-Instruct-v0.2131313https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model in our study, as this model supports a context of up to 32K tokens. However, running this model on ELITR-Bench led to a GPU out-of-memory error on the A100, and thus we discarded it.
@@ A.2 Configuration search on ELITR-Bench-QA’s dev set
The cost of the two-step configuration search amounted to approximately \$150.161616We assessed the cost of performing the evaluation of a single model on the 141 dev set questions to \$3 approximately. As we evaluated 7 models on 6 configurations in the first step, and 7 models on 1 configuration in the second step, this yields \$147. To limit excessive expenses, we used the same model configuration for the different settings we experimented in (single-turn ELITR-Bench-QA, multi-turn ELITR-Bench-QA, and multi-turn ELITR-Bench-Conv).
@@ B.2 Additional results on question type and answer position
B.3 Updated results with recent long-context LLMs
In this section, we provide updated results to include some of the latest long-context LLMs that have been released after the core experiments of this paper have been conducted. In particular, we added the three following models: OpenAI’s GPT-4o, [^17] https://platform.openai.com/docs/models/gpt-4o GradientAI’s LLaMA-3-8B-Instruct-262K [^18] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k which was fine-tuned from the original LLaMA-3 model, [^19] https://llama.meta.com/llama3/ and Microsoft’s Phi-3-Small-128K-Instruct. [^20] https://huggingface.co/microsoft/Phi-3-small-128k-instruct For GPT-4o, we used the same hyperparameters as for GPT-4 and GPT-4.5: nucleus sampling with a temperature of 0.6 and top-p of 0.9. For LLaMA-3-8B-Instruct-262K and Phi-3-Small-128K-Instruct, inference was done by greedy decoding and we used their original chat templates. QA markers and repetition penalty have not been included.
The comparison of these additional models and previously used approaches is provided in Table 11. These results show that GPT-4o performed essentially on par with GPT-4, with a small gain on the dev set. Interestingly, LLaMA-3-8B-Instruct-262K remains close in performance to LLaMA-2 models and only slightly outperforms the best among these, while Phi-3-Small-128K-Instruct beats both LLaMA-2 and LLaMA-3 models by a good margin. It even outperforms GPT-3.5 despite having only 7B parameters. This suggests that this Phi-3 model is a good open-source alternative to more powerful models such as GPT-4 and GPT-4o for our meeting assistant task.
Table 11: Results on single-turn ELITR-Bench-QA updated with recent long-context LLMs: GPT-4o, LLaMA-3-8B-Instruct-262K, and Phi-3-Small-128K-Instruct. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained by a GPT-4 evaluator, on a single seeded run for the dev set and 3 seeded runs for the test set. Boldface numbers correspond to the best performance among proprietary or open-source models.
Model Single-turn ELITR-Bench-QA
Dev Test Mean
GPT-3.5 7.04 7.44 7.24
GPT-4 8.21 8.39 8.30
GPT-4o 8.50 8.44 8.47
LongAlpaca-7B 5.89 5.60 5.75
LongAlpaca-13B 6.17 6.25 6.21
LongChat-7B-v1.5 6.60 5.78 6.19
Vicuna-7B-v1.5 5.42 5.61 5.51
Vicuna-13B-v1.5 5.92 6.52 6.22
LongAlign-7B 6.11 6.46 6.28
LongAlign-13B 6.27 6.33 6.30
LLaMA-3-8B-Instruct-262K 6.83 6.51 6.67
Phi-3-Small-128K-Instruct 7.37 7.34 7.36